OpenStudy (anonymous):

[2.03] Solve: 5(x − 1) = 3(x + 2) − 7 4 −2 −3 2

OpenStudy (anonymous):

2

OpenStudy (anonymous):

5x-5=3x+6-108 2x=-97 x=-97/2